# RAG-Enhanced Commit Message Generation

### v2

@@ 4.3. Model Selection
Table 2 shows the selected CLMs in this paper. Specifically, in the experiment, we adopt the bimodal version for PLMs to acquire the capacities of both natural language and code. For ChatGPT, we use gpt-3.5-turbo API version. For Llama 3 and Gemma, we adopt the instruct version so that LLMs can follow the prompt instructions.
@@ 4.4. Compared Baselines
Hybrid methods: We consider two hybrid methods that take advantage of both IR and NMT techniques, CoRec (Wang et al., 2022), RACE (Shi et al., 2022).
@@ 4.6. Implementation Details
The PLMs used in this study are implemented from the official HuggingFace repository [^5] https://huggingface.co/Salesforce/codet5-base [^6] https://huggingface.co/Salesforce/codet5p-220m-bimodal [^7] https://huggingface.co/uclanlp/plbart-base [^8] https://huggingface.co/microsoft/unixcoder-base, Gemma and Llama 3 are based on the official released model weights and use FP16 precision. ChatGPT used the gpt-3.5-turbo API as of the experiment date, May 1. The fine-tuning of the PLMs was completed on an NVIDIA RTX 4090, utilizing the Adam optimizer to minimize the cross-entropy loss function. The models were trained for 20 epochs using a batch size of 32 and a learning rate of 5e-5, and the same hyperparameters were applied to all four PLMs. Due to the vast size of the source database, which contains millions of entries, we employed the industrial-grade search engine Apache PyLucene (luc, 1999) to calculate BM25 scores. This ensures that the retrieval average time for one query remains acceptably short.
5. Results & Analysis
In this section, we present and discuss the experimental results to answer the proposed four RQs.
5.1. RQ1: Efficacy of Directly Applying CLMs to the CMG Task
At the beginning, we evaluated the CLMs selected for the experiments by directly applying them to the CMG task to assess their feasibility. These CLMs have already been widely applied to various code-related tasks, such as code completion, code summarization, and automated program repair. Additionally, some recent works (Zhang et al., 2024c; Xue et al., 2024; Lopes et al., 2024) have explored the performance of some CLMs, e.g., CodeBERT, Llama, and ChatGPT, on the CMG task. However, the evaluations of these studies were conducted on small sampled datasets, and their model selections were neither comprehensive coverage nor cutting-edge. Therefore, it is necessary to comprehensively evaluate and compare the performance of selected CLMs directly applied to the CMG task before investigating the remaining RQs.
@@ 5.1.2. Analysis
5.2. RQ2: Effectiveness of REACT When Integrating with PLMs
@@
ChatGPT directly 2.50 18.07 17.61
REACT 3.11 $\uparrow$2418.09 $\sim$019.25 $\uparrow$9.3%
@@ 5.3.2. Analysis
REACT can enhance the performance of ChatGPT and Llama 3 in CMG. However, the results for Gemma are somewhat anomalous. Although integrating Gemma into REACT increases its BLEU score by 51%, it leads to a decline in the scores for the other two metrics. This is partly because the focus of evaluation differs among the three metrics. BLEU primarily measures the overlap between generated and reference texts, emphasizing precision. The other two metrics might assess aspects like relevance, fluency, or diversity, which could be adversely affected by the changes that improve BLEU scores. On the other hand, Gemma has the smallest parameter size among the three and has relatively weaker capabilities (according to the EvalPlus Leaderboard [^9] https://evalplus.github.io/leaderboard.html). The enhancement brought by REACT relies on the LLM’s in-context learning ability, and Gemma’s limited capabilities might prevent it from reliably benefiting from the exemplar.
@@
  • Integrating LLMs into the REACT framework significantly enhances their CMG performance.
@@ 8. Conclusions
In this paper, we introduced REACT, a novel retrieval-augmented framework designed to enhance commit message generation by effectively integrating advanced retrieval techniques with various language models with code capabilities. Through comprehensive evaluations, we demonstrated that REACT significantly improves the performance of both pre-trained language models and large language models on the commit message generation task. Experimental results indicate that when CodeT5 is integrated into REACT, its metric scores outperform all baselines and achieve the new SOTA. These findings underscore the efficacy of leveraging retrieval-augmented generation in conjunction with the rich prior knowledge embedded in CLMs, providing a robust and efficient solution for automated commit message generation.
For future work, we will consider incorporating additional information beyond retrieving similar exemplars to enhance the input. For instance, for issue-related commits, we can include the issue’s description and discussion information to aid in generating commit messages. Meanwhile, the RAG paradigm can be extended to more tasks of software engineering.

### v3

@@ 4.3. Model Selection
Table 2 shows the selected CLMs in this paper. Specifically, in the experiment, we adopt the bimodal version for PLMs to acquire the capacities of both natural language and code. For ChatGPT, we use the GPT-4o API version. For Llama 3 and Gemma, we adopt the instruct version so that LLMs can follow the prompt instructions.
@@ 4.4. Compared Baselines
Hybrid methods: We consider two hybrid methods that take advantage of both IR and NMT techniques. They are CoRec (Wang et al., 2022) and RACE (Shi et al., 2022).
@@ 4.6. Implementation Details
The PLMs used in this study are implemented from the official HuggingFace repository [^5] https://huggingface.co/Salesforce/codet5-base [^6] https://huggingface.co/Salesforce/codet5p-220m-bimodal [^7] https://huggingface.co/uclanlp/plbart-base [^8] https://huggingface.co/microsoft/unixcoder-base, Gemma and Llama 3 are based on the official released model weights and use FP16 precision. ChatGPT used the GPT-4o API as of the experiment date, Aug 1. The fine-tuning of the PLMs was completed on an NVIDIA RTX 4090, utilizing the Adam optimizer to minimize the cross-entropy loss function. The models were trained for 20 epochs using a batch size of 32 and a learning rate of 5e-5, and the same hyperparameters were applied to all four PLMs. Due to the vast size of the source database, which contains millions of entries, we employed the industrial-grade search engine Apache PyLucene (luc, 1999) to calculate BM25 scores. This ensures that the retrieval average time for one query remains acceptably short.
5. Results and Analysis
In this section, we present and analysis the experimental results to answer the proposed four RQs.
@@ 5. Results and Analysis
5.1. RQ1: Efficacy of Directly Applying CLMs to the CMG Task
At the beginning, we evaluated the CLMs selected for the experiments by directly applying them to the CMG task to assess their feasibility. These CLMs have already been widely applied to various code-related tasks, such as code completion, code summarization, and automated program repair. Additionally, some recent works (Zhang et al., 2024c; Xue et al., 2024; Lopes et al., 2024) have explored the performance of some CLMs, e.g. CodeBERT, Llama, and ChatGPT, on the CMG task. However, the evaluations of these studies were conducted on small sampled datasets, and their model selections were neither comprehensive coverage nor cutting-edge. Therefore, it is necessary to comprehensively evaluate and compare the performance of selected CLMs directly applied to the CMG task before investigating the remaining RQs.
@@ 5.1.2. Analysis
GPT-4o directly 2.57 20.25 19.03
REACT 3.89 $\uparrow$5120.45 $\sim$019.78 $\uparrow$3.9%
@@
5.2. RQ2: Effectiveness of REACT When Integrating with PLMs
@@ 5.3.2. Analysis
REACT can enhance the performance of ChatGPT and Llama 3 in CMG. However, the results for Gemma are somewhat anomalous. Although integrating Gemma into REACT increases its BLEU score by 51%, it leads to a decline in the scores for the other two metrics. This is partly because the focus of evaluation differs among the three metrics. BLEU primarily measures the overlap between generated and reference texts, emphasizing precision. The other two metrics might assess aspects like relevance, fluency, or diversity, which could be adversely affected by the changes that improve BLEU scores. On the other hand, Gemma has the smallest parameter size among the three and has relatively weaker capabilities (according to the EvalPlus Leaderboard (eva, 2023)). The enhancement brought by REACT relies on the LLM’s in-context learning ability, and Gemma’s limited capabilities might prevent it from effectively following the instructions and also reliably benefiting from the guidance of the exemplar.
@@
  • Integrating LLMs into the REACT framework significantly enhances their CMG performance, especially in terms of BLEU scores.
@@ 5.4. RQ4: Ablation Study of the Hybrid Retriever
5.5. RQ5: Within Project Case Study
Figure 6. A case of with-in project study using GPT-4o integrated with REACT. Under the guidance of the retrieved exemplar, GPT-4o effectively adhered to the commit message writing conventions of the project, achieving a high degree of similarity with the ground truth.
The dataset used in this study was collected from multiple repositories. In fact, most of the existing widely-used CMG datasets are cross-project that do not differentiate between projects, meaning that the diff-message pairs retrieved in REACT may come from other projects. However, in general development practices, commit messages within a specific repository often follow similar conventions. When a developer writes a commit message for a project, they typically refer to the historical commit log to adopt a consistent style. In some cases, the repository’s code of conduct explicitly specifies the commit message conventions [^9] Conventional Commits, https://www.conventionalcommits.org/ that should be followed. The REACT method proposed in this paper, a retrieval-augmented generation approach, intuitively proves to be more practical in this scenario since it can adapt the consistent writing style by following the guidance of retrieved exemplar. In this section, we selected an open-source project for a within-project case study, using experiments to further demonstrate the superiority of REACT in this scenario.
5.5.1. Project Selection
We selected an open-source project from GitHub as the case, Electron [^10] https://github.com/electron/electron, visiting date: Sep 1, 2024., which meets the following criteria: it is highly popular, with 114k stars and 15.3k forks; it has a rich history of commits, exceeding 28,000 records; and it follows modern project organization and open-source contribution processes, with a comprehensive contribution guide and code of conduct. As a result, the quality of commit message writing is notably high.
5.5.2. Dataset Collecting
We scraped the historical commit records from the main branch as of January 1, 2020, and then filtered out commits with overly long code diffs and commit messages (¿1000 tokens). Additionally, we excluded commits authored by bots, such as those by dependabot [^11] https://github.com/dependabot, which typically involve bumping a dependency package to a newer version. In the end, we obtained a dataset containing 3,604 diff-message pairs for the single project Electron.
5.5.3. Results & Analysis
Table 6. Results of the with-in project experiments for Electron
Approach BLEU Rouge-L METEOR
GPT4o 3.41 23.79 13.03
REACT(GPT4o) 10.46 30.42 30.57
We conducted a within-project experiment using GPT-4o. During the retrieval phase of the REACT method, we restricted the retrieval to similar examples only from the current project. As shown in Table 6, the experimental results indicate that in the within-project application scenario, the enhancement to the three metric scores provided by REACT is highly significant. It is able to generate commit messages that closely align with the project’s writing style, gaining higher metric scores than directly generating messages without REACT. Figure 6 shows that GPT-4o, based on the retrieved exemplar, learned the commit message writing convention of this project as “[action]: [description]”. The commit message generated by REACT successfully emulated this style. This further highlights the practical significance of REACT in real-world within-project scenarios, demonstrating that it not only significantly improves metric scores of CMG but also produces more usable commit messages that can be effortlessly adopted by developers.
Key Findings of RQ5:
  • REACT enhances commit message generation in within-project scenarios by adapting to project-specific conventions through examples retrieved from the current project, making the messages more usable and aligned with project standards.
@@ 8. Conclusions
This paper proposed REACT, a retrieval-augmented framework designed to enhance commit message generation by effectively integrating advanced retrieval techniques with various language models with code capabilities. Through comprehensive evaluations, we demonstrated that REACT significantly improves the performance of both pre-trained language models and large language models on the commit message generation task. Experimental results indicate that when CodeT5 is integrated into REACT, its metric scores outperform all baselines and achieve the new SOTA. These findings underscore the efficacy of leveraging retrieval-augmented generation in conjunction with the rich prior knowledge embedded in CLMs, providing a robust and efficient solution for automated commit message generation.
For future work, we will consider incorporating additional information beyond retrieving similar exemplars to enhance the input. For instance, for issue-related commits, we can include the issue’s description and discussion information to aid in generating commit messages. Meanwhile, the RAG paradigm can be extended to more software engineering tasks.
Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant Nos. 62032016 and 62172311.